OpenStudy (anonymous):

Find P(A or B), if events A and B are overlapping. P(A) = 9/18 P(B) = 5/18 P(A and B) = 3/18

OpenStudy (anonymous):

9+5 = 14 - 3 = 11 11/18

OpenStudy (anonymous):

just 9+ = 14 -3=11